how would i find the slope of the line for a problem like this one
-4x-2y=-10
this problem really confuses me while looking for the slope i get lost in the numbers

Answer by venugopalramana(3286)

-4x-2y=-10 ...the slope(m)and intercept(c) form of a st.line is y=mx+c...so transform the equation into that form
-2y=4x-10
y=(-4/2)x-(10/-2)
y=-2x+5..hence comparing with y=mx+c..we get m=slope=-2 and c=intercept=5